# Exploring the Impact of Personality Traits on LLM Bias and Toxicity

### v2

@@ Abstract
With the different roles that AI is expected to play in human life, imbuing large language models (LLMs) with different personalities has attracted increasing research interests. While the “personification” enhances human experiences of interactivity and adaptability of LLMs, it gives rise to critical concerns about content safety, particularly regarding bias, sentiment and toxicity of LLM generation. This study explores how assigning different personality traits to LLMs affects the toxicity and biases of their outputs. Leveraging the widely accepted HEXACO personality framework developed in social psychology, we design experimentally sound prompts to test three LLMs’ performance on three toxic and bias benchmarks. The findings demonstrate the sensitivity of all three models to HEXACO personality traits and, more importantly, a consistent variation in the biases, negative sentiment and toxicity of their output. In particular, adjusting the levels of several personality traits can effectively reduce bias and toxicity in model performance, similar to humans’ correlations between personality traits and toxic behaviors. The findings highlight the additional need to examine content safety besides the efficiency of training or fine-tuning methods for LLM personification. They also suggest a potential for the adjustment of personalities to be a simple and low-cost method to conduct controlled text generation.
@@
Shuo Wang1thanks: Equal contribution.and Renhao Li1,211footnotemark: 1thanks: Under the Joint Ph.D. Program between UM and SIAT.and Xi Chen311footnotemark: 1and Derek F. Wong1thanks: Corresponding author.and Yulin Yuan4and Min Yang233footnotemark: 3 1 NLP2CT Lab, Department of Computer and Information Science, University of Macau 2 Shenzhen Key Laboratory for High Performance Data Mining, Shenzhen Institutes of Advanced Technology, Chinese Academy of Sciences 3 Linguistics and Multilingual Studies, Nanyang Technological University 4 Department of Chinese Language and Literature, University of Macau nlp2ct.{shuo,renhao}@gmail.com, {derekfw,yulinyuan}@um.edu.mo zoexi.chen@ntu.edu.sg, min.yang@siat.ac.cn
@@ 1 Introduction
As the demand for large language models (LLMs) to serve diversified roles continues to grow, the topic of LLM personification has surged in LLM research and development Chen et al. (2024). By simulating specific roles with certain personalities, such as a caring AI friend, LLMs enhance both the task effectiveness and naturalness of human-machine interaction, while providing human-centered problem-solving and enriching interactive experiences Wen et al. (2024). However, one fundamental question remains underexplored in the development of anthropomorphic LLM, that is, the potential toxic language and social biases that different personalities may bring about in the process of personification.
@@
It is well known that LLM generation is not bias-free. In fact, previous studies have evidenced that LLMs not only generate but also amplify social biases Gallegos et al. (2024). Especially, when LLMs are assigned specific identities, they may become even targeted at certain protected characteristics, e.g., gender, race, and a combination of them Chen et al. (2024). While a few studies have paid attention to the toxicity and biases encoded by LLM output during their role plays Zhao et al. (2024), how specific personality traits influence model bias and toxicity has scarcely been examined. This study aims to fill the gap by exploring the biases and toxicity arising from different LLM personalities.
Specifically, we leverage sophisticated personality frameworks developed in social psychology to design theoretically and experimentally sound prompts for LLMs. Previous studies have adopted Big Five and MBTI – two well-known personality tests – to examine LLM performance in general Rao et al. (2023); Frisch and Giulianelli (2024). Being aware that MBTI has long been criticized in psychology research Pittenger (2005); McCrae and Costa Jr (1989), we choose the HEXACO model [^1] https://hexaco.org/, that is further developed from Big Five and that have provided of the-state-of-the-art explanations for moral and behavioral characteristics in psychological studies Pringle et al. (2024). The HEXACO model defines six personality traits as shown in Figure 1. For each dimension, scores range from 0 to 5. In our experiment, a high score for a given personality trait is defined as $\geq 4$, while a low score is defined as $\leq 2$. Leveraging the performance descriptions associated with high and low scores on personality tests, we further design instructions to activate specific personality traits in LLMs. Figure 1 illustrates the HEXACO personality dimensions and outlines the primary evaluation workflow.
@@
To examine the relationships between HEXACO personalities and LLMs’ bias and toxicity output, we employ three relevant datasets, including BOLD Dhamala et al. (2021), RealToxicityPrompt Gehman et al. (2020), and BBQ Parrish et al. (2022). BOLD and RealToxicityPrompt are used to evaluate the model’s performance in text generation tasks, while BBQ is used for QA tasks. They provide different types of toxic language and social biases that allow us to obtain generalizable insight. We also adopt triangulated evaluation metrics, including social bias, verbal sentiment, and language toxicity, to assess the impact of various personality traits on model-generated content.
The data analysis results reveal that LLMs are sensitive to personalities provided by HEXACO-based prompts. They demonstrate a consistent variation in toxic language and social biases, when being assigned with certain personality traits. In particular, adjusting the levels of several personality traits, such as Agreeableness, Openness-to-Experience, and Extraversion, can effectively increase/reduce bias and toxicity in model performance, while giving rise to unwanted flattery which is toxic in a different sense.
@@
The contributions of the study are threefold: (i) It highlights the need to re-examine the outcome of LLM training or fine-tuning for personification besides the efficiency of the training methods (e.g., Jiang et al., 2024); (ii) in the meantime, the findings suggest that the adoption of certain personality traits, as part of in-context learning or fine-tuning, might serve to alleviate the toxicity and social biases encoded during the LLM training process; (iii) they also help LLMs interact with users with different personalities and, furthermore, identify potentially risky input.
@@ 2.1 The Role of Personality Traits in Prejudice and Verbal Aggression
Allport et al. (1954) lay the foundation for prejudice research in The Nature of Prejudice, emphasizing the impact of individual beliefs and values on inter-group relations. Social psychological experimental research demonstrates that individual personality traits play a crucial role in the formation of prejudice and the expression of linguistic aggression Buss and Perry (1992); Sibley et al. (2010); Molero Jurado et al. (2018); Zaki et al. (2024); Ekehammar and Akrami (2007). Crawford and Brandt (2019) indicates that among the Big Five personality traits, Agreeableness, Openness and Extraversion show significant negative correlations with prejudice. Similarly, Hu et al. (2022) demonstrate a negative relationship between Agreeableness personality and verbal aggression. Rafienia et al. (2008) shows that positive Extraversion could lead to positive judgment (e.g., probability rating for positive events) and positive interpretation (e.g., writing a positive story).
@@ 2.2 LLM Personification
Research on LLMs in the fields of role-playing and personification has recently gained popularity. Chen et al. (2024) conduct a systematic review on the personification and role-playing of LLMs, proposing a classification of LLM personas: Demographic Personas, Character Personas, and Individualized Personas. Our research focuses on the persona traits of LLMs, which therefore fall under the Demographic Personas. The review summarizes methods for constructing LLM personas, such as (continuous) pre-training, instruction finetuning, reinforcement learning, and contextual learning. Several studies examine the effectiveness of these methods Jiang et al. (2024); Sorokovikova et al. (2024); Wang et al. (2024); Chen et al. (2024); Zhang et al. (2024).
Among the different studies, Zhang et al. (2024) is one of the few that examines content safety and personality. They focus primarily on 7B open-source models and explore the relationship between the MBTI personality types and models’ ability to remain content-safe. In a similar vein, Wan et al. (2023) introduce the concept of ”personalized bias” in dialogue systems, evaluating how LLMs exhibit biases in role plays based on social categories of a role (e.g., “Asian person” or “Yumi”). The finding is corroborated by Zhao et al. (2024) who find that, although role-playing can improve the reasoning capabilities of LLMs, it also introduces potential risks, particularly in generating stereotypical and harmful outputs. While the few studies have contributed invaluable insight into the potential correlations between personality assignment and LLM toxic and/or biased performance, they have either focused on traditional personality types or social categories, the explanatory force of which is rather constrained.
@@ 3.1 Model Settings
Before exploring how personality influences LLM bias and toxicity, we first evaluate whether the model can indeed take on the different personalities prompted by various personality descriptions from the HEXACO framework. Specifically, we design prompts based on performance descriptions corresponding to high and low scores in each personality dimension. We then administer the HEXACO-100-English personality tests Lee and Ashton (2018) on the selected models to evaluate whether they effectively embody the assigned personalities after prompting. Specific personality activation prompts for LLMs are provided in Appendix A.
@@ 3.2 Datasets
Closed-ended Tasks:
For the closed-ended task, we utilize the multi-choice question answering dataset BBQ-Ambiguous Parrish et al. (2022), which covers 11 bias categories (see Appendix B) and consists of 29,246 QAs, each featuring a target bias option. Ambiguous Contexts in BBQ are used to set up the general situation and introduce the two groups related to the questions, assessing the model’s performance when there is insufficient evidence in the context. The correct answer in all ambiguous contexts is the ”UNKNOWN option”. The ambiguous samples of BBQ are more challenging than the disambiguous samplews, which justifies our decision to focus on it. By evaluating selected models on this dataset, we aim to assess their tendency to select biased responses.
Open-ended Tasks:
For the open-ended task, we use two text generation datasets: BOLD Dhamala et al. (2021) and RealToxicityPrompts Gehman et al. (2020). BOLD is an open-ended language generation dataset that provides English text generation prompts for bias benchmarking across five domains. In our experiments, we randomly sample 600 instances from each domain while ensuring an equal number of samples from each subgroup. If the total sample count is not evenly divisible by the number of subgroups, we round to the nearest integer. This approach ensures diverse and balanced subsets for model evaluation, providing a fair representation of bias levels. The RealToxicityPrompts dataset provides sentence-level prompts de- rived from a large corpus of English web text for toxicity testing. We extract the prompts from its challenge subset to ensure a more rigorous assessment. In total, we have 3,014 samples from the BOLD dataset and 1,199 samples from RealToxicityPrompts.
@@ 3.3 Evaluation Methods
For labeled questions in the closed-ended dataset BBQ, we follow Parrish et al. (2022) and adopt the “bias score in ambiguous contexts” to systematically quantify the extent of bias in the model’s answers:
@@
where acc represents the accuracy of the model output on the given questions. $n_{\text{biased\_ans}}$ and $n_{\text{non-unknown\_ans}}$ represent the number of model outputs that reflect the targeted social bias, and the number of model outputs that do not belong to the ”unknown” choice, respectively. A bias score of 0% indicates no bias in the model’s responses, while 100% means all answers reflect the targeted social bias, and -100% indicates all answers are against the targeted bias. We then quantify the correlation by subtracting $S_{\text{bias}}$ obtained from high-score and low-score personality traits.
For the open-ended text generation tasks, we adopt the Sentiment Reasoner (Vader) score $S_{\text{VAD}}$ Hutto and Gilbert (2014) based on Valence Aware Dictionary and the toxicity score $S_{\text{TOX}}$ from a widely-used toxicity classifier (Perspective API  [^2] https://perspectiveapi.com/). Specifically, Vader is a rule-based model for sentiment analysis that calculates sentiment scores using valence-based lexicons and the combination of the lexicons and rules. For each input, it generates a score $S_{\text{VAD}}$ in the range of -1 to 1, where -1 indicates a negative sentiment and 1 indicates a positive sentiment. Following Dhamala et al. (2021), we utilize a threshold of $\geq 0.5$ to classify positive sentiment, and $\leq-0.5$ to classify negative sentiment, against which the proportions of positive $S^{pos}_{\text{VAD}}$ and negative LLM generations $S^{neg}_{\text{VAD}}$ are calculated.
Figure 2: Evaluation results of three selected LLMs on the HEXACO-100-English test. ”High” indicates the model is prompted with a high-score specific personality trait, ”Low” means the model is prompted with a low-score specific personality trait, and ”Base” refers to the model being prompted without personality instructions.
Besides the sentiment analysis, toxicity scores $S_{\text{TOX}}$ are obtained using an automated toxic language detection tool, Perspective API. The scores represent the probability of a LLM generation being toxic Gehman et al. (2020). To represent the overall level of toxicity of the LLM performance, we report the average score over all data samples.
The sentiment scores and toxicity scores complement each other to provide fine-grained insight into the data. Especially, toxic texts may not necessarily be sentimentally negative (e.g., faltering being sentimentally positive but toxic), while non-toxic texts may not always be sentimentally positive (e.g., expressions of sadness). The discrepancies between the two scores reveal many subtle and complex manifestations of bias and toxicity. Besides checking the two types of scores seperately, we also combine the proportions of positive and negative sentiment classifications $S_{\text{VAD}}$, and toxicity scores $S_{\text{TOX}}$, as both share the same range from 0 to 1:
@@
We then subtract the $S_{\text{open}}$ obtained from high-score and low-score personality traits to quantify the impact, similar to what we did with the closed-ended dataset.
@@
GPT-4o-mini Openness to Experiencelow 0.87 3.73 0.81 4.16 -1.02 -0.15 7.83 -8.01 1.39 1.08 -0.70 0.91
@@
Llama-3.1-70B-instruct Openness to Experiencelow -0.43 3.73 2.05 8.96 -0.13 1.92 8.83 -7.05 2.78 2.12 -2.29 1.86
@@
Qwen2.5-72B-instruct Openness to Experiencelow -1.58 5.66 -0.04 2.66 0.00 0.03 1.67 -6.91 -0.93 0.01 -0.70 -0.01

### v3

@@ Abstract
With the different roles that AI is expected to play in human life, imbuing large language models (LLMs) with different personalities has attracted increasing research interest. While the “personification” enhances human experiences of interactivity and adaptability of LLMs, it gives rise to critical concerns about content safety, particularly regarding bias, sentiment, and toxicity of LLM generation. This study explores how assigning different personality traits to LLMs affects the toxicity and biases of their outputs. Leveraging the widely accepted HEXACO personality framework developed in social psychology, we design experimentally sound prompts to test three LLMs’ performance on three toxic and bias benchmarks. The findings demonstrate the sensitivity of all three models to HEXACO personality traits and, more importantly, a consistent variation in the biases, negative sentiment, and toxicity of their output. In particular, adjusting the levels of several personality traits can effectively reduce bias and toxicity in model performance, similar to humans’ correlations between personality traits and toxic behaviors. The findings highlight the additional need to examine content safety besides the efficiency of training or fine-tuning methods for LLM personification, they also suggest a potential for the adjustment of personalities to be a simple and low-cost method to conduct controlled text generation.
@@
Shuo Wang1thanks: Equal contribution.and Renhao Li1,211footnotemark: 1thanks: Under the Joint Ph.D. Program between UM and SIAT.and Xi Chen311footnotemark: 1and Yulin Yuan4and Min Yang2thanks: Corresponding author.and Derek F. Wong133footnotemark: 3 1 NLP2CT Lab, Department of Computer and Information Science, University of Macau 2 Shenzhen Key Laboratory for High Performance Data Mining, Shenzhen Institutes of Advanced Technology, Chinese Academy of Sciences 3 Linguistics and Multilingual Studies, Nanyang Technological University 4 Department of Chinese Language and Literature, University of Macau nlp2ct.{shuo,renhao}@gmail.com, {derekfw,yulinyuan}@um.edu.mo zoexi.chen@ntu.edu.sg, min.yang@siat.ac.cn
@@ 1 Introduction
With the increasing demand for large language models (LLMs) to serve diversified roles, LLM personification has surged in LLM research and development Chen et al. (2024). By simulating specific roles with certain personalities, such as a caring AI friend, LLMs enhance both the task effectiveness and naturalness of human-machine interaction, while providing human-centered problem-solving and enriching interactive experiences Wen et al. (2024). However, one fundamental question remains underexplored in the development of anthropomorphic LLM, that is, the potential toxic language and social biases that different personalities may bring about in the process of personification.
@@
It is well known that LLM generation is not bias-free. In fact, previous studies have evidenced that LLMs not only generate but also amplify social biases Gallegos et al. (2024). Especially, when LLMs are assigned specific identities, they may become even targeted at certain protected characteristics, e.g., gender, race, and a combination of them Chen et al. (2024). While a few studies pay attention to the toxicity and biases encoded by LLM output during their role plays Zhao et al. (2024), how specific personality traits influence model bias and toxicity has scarcely been examined. This study aims to fill the gap by exploring the biases and toxicity arising from different LLM personalities.
We leverage advanced personality frameworks from social psychology to design theoretically grounded prompts for LLMs. Although previous work has used popular models like the Big Five and MBTI to evaluate LLM behavior Rao et al. (2023); Frisch and Giulianelli (2024), MBTI has been widely criticized for its low reliability, due to its rigid dichotomization of personality traits and poor test-retest consistency—nearly 50% of individuals change types over time Matz et al. (2016); Howes and Carskadon (1979).In contrast, the HEXACO model builds on the Big Five by adding a sixth dimension, honesty-humility, which has proven valuable in predicting morally relevant behaviors such as cheating, free-riding, ethical leadership, short-term mating, and gambling Lee and Ashton (2020). Although some researchers argue that honesty-humility can operate independently of other personality models Howard and Van Zandt (2020), recent evidence shows that HEXACO outperforms the Big Five in explaining health-related behaviors, largely due to the unique variance contributed by honesty-humility Pletzer et al. (2024). Given these advantages and the growing critique of MBTI in psychological research Pittenger (2005); McCrae and Costa Jr (1989), we adopt the HEXACO model [^1] https://hexaco.org/ as the basis for our experimental design. HEXACO defines six personality dimensions (Figure 1), each scored from 0 to 5. In our experiments, scores $\geq$ 4 are considered high, and scores $\leq$ 2 are low. Based on the descriptive behaviors associated with these high and low scores, we design targeted instructions to activate specific personality traits in LLMs. Figure 1 shows the HEXACO dimensions and the main evaluation workflow.
@@
To examine the relationships between HEXACO personalities and LLM bias and toxicity output, we employ three relevant datasets, including BOLD Dhamala et al. (2021), RealToxicityPrompt Gehman et al. (2020), and BBQ Parrish et al. (2022). The first two datasets assess model performance in text generation tasks, while the third evaluates quality control in bias detection. Together, they provide diverse forms of toxic language and social biases, enabling robust and generalizable insights. We also adopt triangulated evaluation metrics, including social bias, verbal sentiment, and language toxicity, to assess the impact of various personality traits on model-generated content. Our analysis reveals that LLMs are sensitive to personalities provided by HEXACO-based prompts. They demonstrate a consistent variation in toxic language and social biases when assigned certain personality traits. In particular, adjusting the levels of several personality traits, such as Agreeableness, Openness-to-Experience, and Extraversion, can effectively increase/reduce bias and toxicity in model performance, while giving rise to unwanted flattery that is toxic in a different sense.
@@
The contributions of this study are threefold: (i) It highlights the need to re-examine the outcome of LLM training for personification, besides the effectiveness of training methods; (ii)the findings also suggest that the adoption of certain personality traits, as part of in-context learning, might serve to alleviate the toxicity and biases of LLM generation; (iii) they also help LLMs interact with users with diverse personalities and further identify potentially risky input.
@@ 2.1 The Role of Personality Traits in Prejudice and Verbal Aggression
Allport et al. (1954) lay the foundation for prejudice research in The Nature of Prejudice, emphasizing the impact of individual beliefs and values on inter-group relations. Social psychological experimental research demonstrates that individual personality traits play a crucial role in the formation of prejudice and the expression of linguistic aggression Buss and Perry (1992); Sibley et al. (2010); Molero Jurado et al. (2018); Zaki et al. (2024); Ekehammar and Akrami (2007). Crawford and Brandt (2019) indicates that among the Big Five personality traits, Agreeableness, Openness, and Extraversion show significant negative correlations with prejudice. Similarly, Hu et al. (2022) demonstrate a negative relationship between Agreeableness personality and verbal aggression. Rafienia et al. (2008) show that positive Extraversion could lead to positive judgment and interpretation.
@@ 2.2 LLM Personification
Research on LLMs in the fields of role-playing and personification has recently gained popularity. Chen et al. (2024) conduct a systematic review on the personification and role-playing of LLMs, proposing a classification of LLM personas: Demographic Personas, Character Personas, and Individualized Personas. Our research focuses on the persona traits of LLMs, which therefore fall under the Demographic Personas. The review summarizes methods for constructing LLM personas, such as pre-training, instruction fine-tuning, reinforcement learning, and contextual learning. Several studies examine the effectiveness of these methods Jiang et al. (2024); Sorokovikova et al. (2024); Wang et al. (2024); Chen et al. (2024). Among these studies, Zhang et al. (2024) is one of the few that examines content safety and personality. They focus primarily on 7B open-source models and explore the relationship between MBTI personality types and model safety. In a similar vein, Wan et al. (2023) introduce the concept of “personalized bias” in dialogue systems, evaluating how LLMs exhibit biases in role plays based on social categories of a role. The finding is corroborated by Zhao et al. (2024), who find that although role-playing can improve the reasoning capabilities of LLM, it also introduces potential risks, particularly in generating stereotypical and harmful outputs. While the few studies have contributed invaluable insight into the potential correlations between personality assignment and LLM toxic and/or biased performance, they have either focused on traditional personality types or social categories, the explanatory force of which is rather constrained.
@@ 3.1 Model Settings
Before exploring how personality influences LLM bias and toxicity, we first evaluate whether the model can indeed take on the different personalities prompted by various personality descriptions from the HEXACO framework. Specifically, we design prompts based on performance descriptions corresponding to high and low scores in each personality dimension. We then administer the HEXACO-100-English personality tests Lee and Ashton (2018) on the selected models to evaluate whether they effectively embody the assigned personalities after prompting. Specific personality activation prompts are provided in Appendix A.
@@ 3.2 Datasets
For the closed-ended task, we utilize the multi-choice question answering dataset BBQ-Ambiguous Parrish et al. (2022), which covers 11 bias categories (see Appendix B) and consists of 29,246 QAs, each featuring a target bias option. Ambiguous Contexts in BBQ are used to set up the general situation and introduce the two groups related to the questions, assessing the model’s performance when there is insufficient evidence in the context. The correct answer in all ambiguous contexts is the “UNKNOWN option”. The ambiguous samples of BBQ are more challenging than the disambiguous samples, which justifies our decision to focus on it. By evaluating selected models on this dataset, we aim to assess their tendency to select biased responses.
For the open-ended task, we use two text generation datasets: BOLD Dhamala et al. (2021) and RealToxicityPrompts Gehman et al. (2020). BOLD is an open-ended language generation dataset that provides English text generation prompts for bias benchmarking across five domains. In our experiments, we randomly sample 600 instances from each domain while ensuring an equal number of samples from each subgroup. If the total sample count is not evenly divisible by the number of subgroups, we round to the nearest integer. This approach ensures diverse and balanced subsets for model evaluation, providing a fair representation of bias levels. The RealToxicityPrompts dataset provides sentence-level prompts de- derived from a large corpus of English web text for toxicity testing. We extract the prompts from its challenge subset to ensure a more rigorous assessment. In total, we have 3,014 samples from the BOLD dataset and 1,199 samples from RealToxicityPrompts.
@@ 3.3 Evaluation Methods
For labeled questions in the closed-ended dataset BBQ, we follow Parrish et al. (2022) and adopt the “bias score in ambiguous contexts” to quantify the extent of bias in the model’s answers:
@@
where acc is the accuracy of the model output on the given questions. $n_{\text{biased\_ans}}$ and $n_{\text{non-unknown\_ans}}$ represent the number of model outputs that reflect the targeted social bias, and the number of model outputs that do not belong to the ”unknown” choice, respectively. A bias score of 0% indicates that there is no bias in the responses of the model, while $100\%$ means that all answers reflect the targeted social bias, and $100\%$ indicates that all responses are against the targeted bias. We then quantify the correlation by subtracting $S_{\text{bias}}$ obtained from high-score and low-score personality traits.
For the open-ended text generation tasks, we adopt the Sentiment Reasoner (Vader) score $S_{\text{VAD}}$ Hutto and Gilbert (2014) based on the Valence Aware Dictionary and the toxicity score $S_{\text{TOX}}$ from a widely used toxicity classifier (Perspective API  [^2] https://perspectiveapi.com/). Specifically, Vader is a rule-based model for sentiment analysis that calculates sentiment scores using valence-based lexicons and the combination of the lexicons and rules. For each input, it generates a score $S_{\text{VAD}}$ in the range of -1 to 1, where -1 indicates a negative sentiment and 1 indicates a positive sentiment. Following Dhamala et al. (2021), we utilize a threshold of $\geq 0.5$ to classify positive sentiment, and $\leq-0.5$ to classify negative sentiment, against which the proportions of positive $S^{pos}_{\text{VAD}}$ and negative LLM generations $S^{neg}_{\text{VAD}}$ are calculated.
Figure 2: Evaluation results of three selected LLMs on the HEXACO-100-English test. “High” indicates the model is prompted with a high-score specific personality trait, “Low” means the model is prompted with a low-score specific personality trait, and “Base” refers to the model being prompted without personality instructions.
In addition to sentiment analysis, the toxicity scores $S_{\text{TOX}}$ are obtained using a toxic language detection tool, Perspective API. The scores represent the probability of an LLM generation being toxic Gehman et al. (2020).
Sentiment scores and toxicity scores complement each other to provide fine-grained insight into the data. Especially, toxic texts may not necessarily be sentimentally negative (e.g., faltering being sentimentally positive but toxic), while non-toxic texts may not always be sentimentally positive (e.g., expressions of sadness). The discrepancies between the two scores reveal many subtle and complex manifestations of bias and toxicity. Besides checking the two types of scores separately, we also combine the proportions of positive and negative sentiment classifications $S_{\text{VAD}}$, and toxicity scores $S_{\text{TOX}}$, as both share the same range from 0 to 1:
@@
We then subtract the $S_{\text{open}}$ obtained from high-score and low-score personality traits to quantify the impact.
@@
GPT-4o-mini
Openness to Experiencelow 0.87 3.73 0.81 4.16 -1.02 -0.15 7.83 -8.01 1.39 1.08 -0.70 0.91
@@
Llama-3.1-70B-instruct
Openness to Experiencelow -0.43 3.73 2.05 8.96 -0.13 1.92 8.83 -7.05 2.78 2.12 -2.29 1.86
@@
Qwen2.5-72B-instruct
Openness to Experiencelow -1.58 5.66 -0.04 2.66 0.00 0.03 1.67 -6.91 -0.93 0.01 -0.70 -0.01
To assess the robustness of our findings, we conduct multiple evaluations on the BBQ dataset using different rewritten prompts (see Appendix D). We also perform two supplementary evaluations—on knowledge QA and summarization tasks—to assess whether our methodology impacts model performance on general tasks (see Appendix E).